# Learning Delays in Spiking Neural Networks using Dilated Convolutions with Learnable Spacings

### v1

@@ 4.3 Ablation study
In this section, we conduct control experiments aimed at assessing the effectiveness of our delay learning method. The model trained using our full method will be referred to as Decreasing $\sigma$, while Constant $\sigma$ will refer to a model where the standard deviation $\sigma$ is constant and equal to the minimum value of $0.5$ throughout the training. Additionally, Fixed random delays will refer to a model where delays are initialiazed randomly and not learned, while only weights are learned. Meanwhile Decreasing $\sigma$ Fixed weights will refer to a model where the weights are fixed and only delays are learned with a decreasing $\sigma$. Finally, No delays denotes a standard SNN without delays. To ensure equal parameter counts across all models (for fair comparison), we increased the number of hidden neurons in the No delays case. Moreover, to make the comparison even fairer, all models have the same initialization for weights, and if required, the same initialization for delays.
@@
Figure 5: Barplots of test accuracies on SHD and SSC datasets for different models. With (a): fully connected layers (FC) and (b): with sparse synaptic connections (S). Reducing the number of synaptic connections of each neuron to ten, for both SHD and SSC.
@@
To explain this, we hypothesized that a random uniformly distributed set of delay positions will most likely cover the whole possible temporal range. This hypothesis is plausible given the fact that the number of synaptic connections vastly outnumbers the total possible discrete delay positions for each kernel. Therefore, as the number of synaptic connections within a layer grows, the necessity of moving delay positions away from their initial state diminishes. And only tuning the weights of this set of fixed delays is enough to achieve comparable performance to delay learning.
In order to validate this hypothesis, we conducted a comparison using the same models with a significantly reduced number of synaptic connections. We applied fixed binary masks to the network’s synaptic weight parameters. Specifically, for each neuron in the network we reduced the number of its synaptic connections to ten for both datasets (except for the No delays model, which has more connections to ensure equal parameters counts). This corresponds to 96% sparsity for SHD and 98% sparsity for SSC. With the number of synaptic connections reduced, it is unlikely that the random uniform initialization of delay positions will cover most of the temporal range. Thus, specific long term dependencies will need to be learned by moving the delays.
The test accuracies corresponding to this control test are shown in Figure 4(b). And it illustrates the difference in performance between the Fixed random delays model and the Decreasing/Constant $\sigma$ models in the sparse case. This enforces our hypothesis and shows the need to perform this control test for delay learning methods. Furthermore, it also indicates the effectiveness of our method.
In addition, we also tested a model where only the delays are learned while the synaptic weights are fixed (Decreasing $\sigma$ Fixed weights). It can be seen that learning only the delays gives acceptable results in the fully connected case (in agreement with beyondweights ) but not in the sparse case. To summarize, it is always preferable to learn both weights and delays (and decreasing $\sigma$ helps). If one has to choose, then learning weights is preferable, especially with sparse connectivity.
@@ 5 Conclusion
In this paper, we propose a method for learning delays in feedforward spiking neural networks using dilated convolutions with learnable spacings (DCLS). Every synaptic connection is modelled as a 1D Gaussian kernel centered on the delay position, and DCLS is used to learn the kernel positions (i.e delays). The standard deviation of the Gaussians is decreased throughout training, such that at the end of training we obtain a SNN model with one discrete delay per synapse, which could potentially be compatible with neuromorphic implementations. We show that our method outperforms the state-of-the-art in both temporal spiking benchmarks SHD and SSC, while using fewer parameters than previous proposals. And we also perform a rigorous control test that demonstrates the effectiveness of our method. Future work could investigate the use of other kernel functions than the Gaussian, or applying our method to other network architectures like convolutional networks.
Limitations
The primary limitations of our work revolve around the compatibility and constraints of our delay learning method. Specifically, our method is limited to offline training conducted in discrete-time simulations, and it cannot handle recurrent connections. Additionally, a maximum delay limit, which corresponds to the size of the kernel, must be predetermined and fixed before the learning process.
Computational resources
This project required about 500 GPU hours on a single Nvidia Tesla T4 GPU with two Intel(R) Xeon(R) CPUs @ 2.20 GHz threads. Given this hardware configuration, a single training session lasted approximately 1 hour for the SHD runs, while for the SSC runs, a single training session lasted around 7 hours. The available computing resources allowed us to perform the required calculations efficiently, leading to accurate and competitive outcomes within a reasonable time.

### v3

@@ 4.3 Ablation study
In this section, we conduct control experiments aimed at assessing the effectiveness of our delay learning method. The model trained using our full method will be referred to as Decreasing $\sigma$ (specifically, we use the 2L-1KC version), while Constant $\sigma$ will refer to a model where the standard deviation $\sigma$ is constant and equal to the minimum value of $0.5$ throughout the training. Additionally, Fixed random delays will refer to a model where delays are initialized randomly and not learned, while only weights are learned. Meanwhile, Decreasing $\sigma$ Fixed weights will refer to a model where the weights are fixed and only delays are learned with a decreasing $\sigma$. Finally, No delays denotes a standard SNN without delays. To ensure equal parameter counts across all models (for fair comparison), we increased the number of hidden neurons in the No delays - wider case, and increased the number of layers instead in the No delays - deeper case. Moreover, to make the comparison even fairer, all models have the same initialization for weights and, if required, the same initialization for delays.
@@
Figure 4: Barplots of test accuracies on SHD and SSC datasets for different models. With (a): fully connected layers (FC) and (b): sparse synaptic connections (S). Reducing the number of synaptic connections of each neuron to ten for both SHD and SSC.
@@
To explain this, we hypothesized that a random uniformly distributed set of delay positions will likely cover the whole temporal range. This hypothesis is plausible given the fact that the number of synaptic connections vastly outnumbers the total possible discrete delay positions for each kernel. Therefore, as the number of synaptic connections within a layer grows, the necessity of moving delay positions away from their initial state diminishes. And only tuning the weights of this set of fixed delays is enough to achieve comparable performance to delay learning.
In order to validate this hypothesis, we conducted a comparison using the same models with a significantly reduced number of synaptic connections. We applied fixed binary masks to the network’s synaptic weight parameters. Specifically, for each neuron in the network, we reduced the number of its synaptic connections to ten for both datasets (except for the No delays model, which has more connections to ensure equal parameter counts). This corresponds to 96% sparsity for SHD and 98% sparsity for SSC. With the number of synaptic connections reduced, it is unlikely that the random uniform initialization of delay positions will cover most of the temporal range. Thus, specific long-term dependencies will need to be learned by moving the delays.
The test accuracies corresponding to this control test are shown in Figure 3(b). It illustrates the difference in performance between the Fixed random delays model and the Decreasing/Constant $\sigma$ models in the sparse case. This enforces our hypothesis and shows the need to perform this control test for delay learning methods. Furthermore, it also indicates the effectiveness of our method.
In addition, we also tested a model where only the delays are learned while the synaptic weights are fixed (Decreasing $\sigma$ Fixed weights). It can be seen that learning only the delays gives acceptable results in the fully connected case (in agreement with Grappolini & Subramoney (2023)) but not in the sparse case. To summarize, it is always preferable to learn both weights and delays (and decreasing $\sigma$ helps). If one has to choose, then learning weights is preferable, especially with sparse connectivity.
@@ 5 Conclusion
In this paper, we propose a method for learning delays in feedforward spiking neural networks using dilated convolutions with learnable spacings (DCLS). Every synaptic connection is modeled as a 1D Gaussian kernel centered on the delay position, and DCLS is used to learn the kernel positions (i.e. delays). The standard deviation of the Gaussians is decreased throughout training, such that at the end of training, we obtain a SNN model with one discrete delay per synapse, which could potentially be compatible with neuromorphic implementations. We show that our method outperforms the state-of-the-art in the temporal spiking benchmarks SHD and SSC and the non-spiking benchmark GSC-35 while using fewer parameters than previous proposals. Finally, we also perform a rigorous control test that demonstrates the effectiveness of our delay learning method. Future work will investigate the use of other kernel functions than the Gaussian or applying our method to other network architectures like convolutional networks.